6x-5y=-67
7x+83=y
7m+n=13
m-9n=11
4x+5y=15
-6x+y=37
6x+9y=-33
-3x+y=22
Im suppose to solve using the substition method, I am confused!

Answer by stanbon(75887)

6x-5y=-67
7x+83=y
---
Substitute for "y" and solve for "x":
6x -5(7x+83) = -67
6x-35x-415 = -67
-29x = 348
x = -12
----
Solve for "y:
y = 7x+83
y = 7*-12 + 83
y = 0
===================================================

7m+n=13
m-9n=11
---
Solve the bottom equation for "m":
m = 9n+11
---
Substitute into the 1st equation and solve for "n":
7(9n+11) + n = 13
63n + 77 + n = 13
64n = - 64
n = -1
----
Substitute for "n" an solve for "m":
m = 8n+11
m = 8*-1 + 11
m = 3
==================
Work the others in a similar manner.
Cheers,
Stan H.
=============



4x+5y=15
-6x+y=37
6x+9y=-33
-3x+y=22
Im suppose to solve using the substition method